OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. -2x + 7y=7

OpenStudy (anonymous):

y=1 x=-7/2

OpenStudy (anonymous):

First solve for y -2x+7y=7 7y=7+2x y=(7/7)+(2/7)x y=1+(2/7)x For the y-intercept, substitute 0 for x y=1+(2/7)*0 y=1 For the x-intercept, substitute 0 for y 0=1+(2/7)x -1=(2/7)x -7=2x -7/2=x